# Control Barrier Functions via Minkowski Operations for Safe Navigation among Polytopic Sets

### v1

@@ IV-A Optimization-Based SDF Formulation in $\mathcal{M}_{\mathcal{D}}$-space
$\begin{split}&\min\,\,\|\bm{z}\|^{2}_{2}\quad\text{subject to}\,\,\,\,\text{subject to}\,\,\,\,A^{\mathcal{C}}\bm{z}\leq b^{\mathcal{C}}\\ &\,\,\text{with}\,\,\,\,\text{dist}(\bm{0}_{c},\mathcal{O}^{\mathcal{C}})=\|\bm{z}^{*}\|_{2},\end{split}$ (10)
@@ IV-C Derivative Calculation Using Differentiable Optimization
Since (10) is a QP in standard form, the Jacobian can be computing using the framework introduced in [41]. We start by finding the Karush-Kuhn-Tucker (KKT) conditions of (10), where the stationarity and complementary slackness conditions can be compactly expressed in matrix form as:
@@
where $I_{d}$ is the $d$-dimensioanl identity matrix, $\bm{\lambda}\in\mathbb{R}^{\ell_{\mathcal{C}}}$ is the dual variable of (10), and $D(\cdot)$ creates a diagonal matrix from a vector. Let $\bm{\xi}:=(\bm{z},\bm{\lambda})$. Thus, a convex optimization solver can be viewed as a root-finding method that seeks the optimal solution $\bm{\xi}^{*}(\bm{x})$ by solving $G(\bm{\xi}^{*})=0$ (suppressing the dependency of $\bm{\xi}$ on $\bm{x}$ for brevity). The partial derivative of $\bm{z}^{*}$ w.r.t. $\bm{x}$ can be computed by differentiating the KKT conditions using the implicit function theorem [44]:
@@
where $d_{\bm{x}}A^{\mathcal{C}}\in\mathbb{R}^{\ell_{\mathcal{C}}\times d\times n}$ is a 3D tensor and $d_{\bm{x}}b^{\mathcal{C}}\in\mathbb{R}^{\ell_{\mathcal{C}}\times n}$ is a matrix, representing how the bounding constraints of CO, $A^{\mathcal{C}},b^{\mathcal{C}}$, changes w.r.t. state $\bm{x}$. Note that $\bm{z}^{*},\bm{\lambda}^{*},A^{\mathcal{C}},b^{\mathcal{C}}$ all depend on the state $\bm{x}$, which we denote compactly as $\bm{z}^{*}(\bm{x}),\bm{\lambda}^{*}(\bm{x}),A^{\mathcal{C}}(\bm{x}),b^{\mathcal{C}}(\bm{x})$. The Jacobian of the optimal solution $\bm{z}^{*}$ can be extracted from the associated block (16), which also provides the derivatives of dual variables. The partial derivative of $h$ can be computed using (15) to 17.
@@ V Case Study and Implementation
Figure 2: Top-Left: The robot’s trajectory (green) moving left-to-right from the start (${t=0}$) to the goal (${t=8}$). Snapshots show the robot (blue) avoiding the obstacle (red). Top-Right: The corresponding motion of the CSO (green) moving right-to-left from ${t=0}$ to ${t=8}$ in $\mathcal{M}_{\mathcal{D}}$-space  (see video https://youtu.be/3Dh0gtDW8bE). Bottom: Control input $\bm{u}(t)$, and CBF $h(\bm{x})$ profiles. Both control inputs satisfy actuator limits and safety constraints.

### v2

@@ IV-A Optimization-Based SDF Formulation in $\mathcal{M}_{\mathcal{D}}$-space
$\begin{split}&\min\,\,\|\bm{z}\|^{2}_{2}\quad\text{subject to}\,\,\,\,A^{\mathcal{C}}\bm{z}\leq b^{\mathcal{C}}\\ &\,\,\text{with}\,\,\,\,\text{dist}(\bm{0}_{c},\mathcal{O}^{\mathcal{C}})=\|\bm{z}^{*}\|_{2},\end{split}$ (10)
@@ IV-C Derivative Calculation Using Differentiable Optimization
Since (10) is a QP in standard form, the Jacobian can be computed using the framework introduced in [41]. We start by finding the Karush-Kuhn-Tucker (KKT) conditions of (10), where the stationarity and complementary slackness conditions can be compactly expressed in matrix form as:
@@
where $I_{d}$ is the $d$-dimensional identity matrix, $\bm{\lambda}\in\mathbb{R}^{\ell_{\mathcal{C}}}$ is the dual variable of (10), and $D(\cdot)$ creates a diagonal matrix from a vector. Let $\bm{\xi}:=(\bm{z},\bm{\lambda})$. Thus, a convex optimization solver can be viewed as a root-finding method that seeks the optimal solution $\bm{\xi}^{*}(\bm{x})$ by solving $G(\bm{\xi}^{*})=0$ (suppressing the dependency of $\bm{\xi}$ on $\bm{x}$ for brevity). The partial derivative of $\bm{z}^{*}$ w.r.t. $\bm{x}$ can be computed by differentiating the KKT conditions using the implicit function theorem [44]:
@@
where $d_{\bm{x}}A^{\mathcal{C}}\in\mathbb{R}^{\ell_{\mathcal{C}}\times d\times n}$ is a 3D tensor and $d_{\bm{x}}b^{\mathcal{C}}\in\mathbb{R}^{\ell_{\mathcal{C}}\times n}$ represents how the bounding constraints of CO, $A^{\mathcal{C}},b^{\mathcal{C}}$, changes w.r.t. state $\bm{x}$. Note that $\bm{z}^{*},\bm{\lambda}^{*},A^{\mathcal{C}},b^{\mathcal{C}}$ all depend on the state $\bm{x}$, which we denote compactly as $\bm{z}^{*}(\bm{x}),\bm{\lambda}^{*}(\bm{x}),A^{\mathcal{C}}(\bm{x}),b^{\mathcal{C}}(\bm{x})$. The Jacobian of the optimal solution $\bm{z}^{*}$ can be extracted from the associated block (16), which also provides the derivatives of dual variables. The partial derivative of $h$ can be computed using (15) to 17.
@@ V Case Study and Implementation
Figure 2: Top-Left: The robot’s trajectory (green) moving left-to-right from the start (${t=0}$) to the goal (${t=8}$). Snapshots show the robot (blue) avoiding the obstacle (red). Top-Right: The corresponding motion of the CO (green) moving right-to-left from ${t=0}$ to ${t=8}$ in $\mathcal{M}_{\mathcal{D}}$-space  (see video https://youtu.be/3Dh0gtDW8bE). Bottom: Control input $\bm{u}(t)$, and CBF $h(\bm{x})$ profiles. Both control inputs satisfy actuator limits and safety constraints